What is the value of KP for the reaction of nitrogen and oxygen to make dinitrogen monoxide if the equilibrium partial pressures of nitrogen is 1.3365 atm, the partial pressure of oxygen is 1.0883 atm and the partial pressure of dinitrogen monoxide is 0.0254 atm?
Answer #1

The balanced equation is:

2 N2(g) + O2(g) <-> 2 N2O(g)

Kp = p(N2O)^2 / p(N2)^2 * p(O2)
= (0.0254)^2 / (1.3365)^2 * 1.0883
= (6.4516*10^-4) / (1.944)
= 3.319*10^-4

Answer: 3.319*10^-4